Unite Inc. maintains a periodic inventory system and uses the gross method to record purchases. The following transactions occurred during the month of March 2020 for its major inventory line.

a. Purchase of merchandise inventory on March 1, 2020, for $14,400 on account, terms 1/10, n/30.

b. Paid $144 cash for freight charges on March 1, 2020, related to the purchase.

c. Returned $108 of merchandise on March 5, 2020, and received a credit from the vendor.

d. Paid the balance due to the vendor on March 8, 2020.

e. Sold merchandise inventory on March 15, 2020, for $9,000.

Prepare journal entries for transactions a through e.

Note: Round your answers to the nearest whole dollar.
Answer #1
Journal Entries
Date Account Title Debit Credit
March 01,2020 Purchases /Cost Of Good Sold $14400
Account Payable 14400
(Inventory purchased)
March 01,2020 Cost Of Good Sold 144
Cash 144
(Frieght paid in cash for buying inventory)
March 05,2020 Account Payable 108
Purchases 108
(purchase return)
March 08,2020 Account Payable 14292
Cash 14149
Cash Discount Received 143
[14400-108]*99% =14149
Discount = 14292-14149 =$143
March 15,2020 Account Receivable / Cash 9000
Sales 9000
(goods sold )
(Note : In the question whether good sold on credit or cash has not given
hence student can assume either cash / account receivable option)